Consider two neighboring island countries called A
Answer #1

Arcadia?s opportunity cost of producing 1 bushel of corn is = 16/8 = 2 pairs of jeans

Felicidad?s opportunity cost of producing 1 bushel of corn is = 20/5 = 4 pairs of jeans

Thus, Arcadia has a comparative advantage in the production of corn, and Felicidad in the production of jeans

Arcadia will use 1+3 = 4 million hours of labor per week to produce corn

Thus, production of corn in Arcadia = 8 * 4 = 32 million bushels per week

Felicidad will use 1+3 = 4 million hours per week to produce jeans

Thus, production of jeans in Felicidad = 20*4 = 80 million pairs per week

Because of specialization, the total production of corn has increased by 32-23 = 9 million

And production of jeans has increased by 80-68 = 12 million pairs

Arcadia Felicidad
Corn Jeans Corn Jeans
Without trade
Production 8 48 15 20
Consumption 8 48 15 20
With trade
Prodcution 32 0 0 80
Trade Action -18 54 18 -54
Consumption 14 54 18 26
Gains from Trade
Increase in consumption 6 6 3 6